# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

### v1

@@ 7. Conclusion
LLM-based APR has progressed from early prototypes to a diverse ecosystem that includes fine-tuned models, prompt-centric methods, scripted pipelines, and autonomous agents. Each paradigm balances accuracy, cost, and engineering effort in a distinct manner, and no single approach prevails across all repair scenarios. Future gains are more likely to come from disciplined evaluation, smarter context selection, deeper integration with existing developer workflows, and stronger safeguards that keep patches correct, secure, and maintainable.
Comprehensive Overview.
Table 3 consolidates 63 representative LLM-based APR systems, aligning each with its paradigm, training corpus, base model, benchmark, and headline result. This overview enables readers to see how design choices impact repair performance easily and sets the stage for the forward-looking discussion that follows.
LLM-based APR: A Turning Point.
Program repair has moved beyond single-shot prompts that fix simple bugs. Modern agentic systems can scan real repositories, interpret failing tests, and submit production-ready patches. These systems already significantly surpass traditional techniques on established benchmarks. They also open the door to repairing cross-file, multilingual, and security-critical defects that prior methods rarely addressed.
LLM-based APR: A Path Forward.
Turning promising prototypes into dependable development infrastructure will require several advances. Benchmarks must be free of data leakage and patch over-fitting; training pipelines must respect practical resource limits; and retrieval engines must reliably surface the most relevant context. Agentic planners need to reason efficiently and stay within predictable time and money costs. Meeting these challenges will elevate LLM-based APR from an emerging research topic to a routine part of modern software engineering.

### v2

@@ 12. Conclusion
LLM-based software repair has evolved from early single-prompt experiments to a set of distinct paradigms that combine different forms of control over the repair loop with different degrees of model adaptation. This survey provides a structured view of the field by proposing a taxonomy based on control authority and parameter adaptation, with retrieval and analysis augmentation as orthogonal layers; mapping 62 representative systems into this framework; and consolidating their benchmarks, evaluation assumptions, and pass@k definitions into a common set of tables.
Taken together, the taxonomy, corpus, and evaluation summary clarify the tradeoffs among fine-tuning, prompting, procedural pipelines, and agentic frameworks, and show how design choices and evaluation protocols jointly shape reported repair success. The synthesis points to the most important directions for future work: stronger, less contaminated benchmarks; more transparent, comparable reporting of pass@k and resource budgets; more selective, repository-aware retrieval and analysis pipelines; and cost-aware planning for multi-step workflows. Addressing these issues is central to turning LLM-based repair from promising prototypes into dependable infrastructure in everyday software engineering practice.